OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y <= x - 5 y >= -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (anonymous):

i do not know how to solve these type of equations.

OpenStudy (anonymous):

5, -2

OpenStudy (anonymous):

but how do you solve it and thank you

OpenStudy (anonymous):

we have y≤x−5 and y≥−x−4 let's substitute each point and check a)5,2 in y<=x-5 2<=5-5 => 2<=0 this is not true so this is not the answer b) 5,-2 in -2<=5-5 -2<=5-5 -2<=0 this is true let's check with the other inequality y>=-x-4 -2>=-5-4 true so option B is one of the answers c) (-5,-2) y<=x-5 -2<=-5-5 -2<=-10 false d)(-5,2) 2<=-5-5 2<=-10 this is also not true SO the answer is option B

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

next time just type it up on google.